OpenStudy (anonymous):

HELP! Medal and fan to the best help:) Let f(x) = 9x - 2 and g(x) = -x + 3. Find f(g(x)). -9x - 2 -9x + 5 -9x + 25 -9x + 27

OpenStudy (ivettef365):

just substitute the x on function f(x) for (-x+3)

OpenStudy (anonymous):

Alright what do I do next?

OpenStudy (ivettef365):

well you have f(x) = 9x -2 9(-x+3) -2 = -9x + 27 -2 = -9x + 25

OpenStudy (anonymous):

That makes so much sense!! Think that you can help me with another few?

OpenStudy (ivettef365):

sure

OpenStudy (anonymous):

Let f(x) = -4x + 7 and g(x) = 10x - 6, find f(g(x)). -40x + 24 -40x + 31 -40x + 64 -40x + 70

OpenStudy (ivettef365):

same process as previews problem substitute the x from f(x) with (10x-6)

OpenStudy (anonymous):

-40x + 64?

OpenStudy (ivettef365):

how did you get to +64 ??

OpenStudy (ivettef365):

you have f(x) = -4x + 7 = -4(10x -6)+7 = -40x +24 + 7 = -40x + 31

OpenStudy (anonymous):

Why would we not use that equation (-4x + 7) on the other side (g(x) = 10x - 6)?

OpenStudy (ivettef365):

the question is f(g{x}), that means you are inserting g(x) inside the function f(x)

OpenStudy (ivettef365):

if the question had been g(f{x}), then you insert f(x) inside g(x)

OpenStudy (anonymous):

Oh okay, this is all making sense now, its all about reading the question properly! Let f(x) = 12/4x + 2 Find f(-1). -6 -2 2 6 How would I go about solving this?

OpenStudy (anonymous):

@ivettef365 ?

OpenStudy (ivettef365):

on his one just substitute the x from f(x) for -1

OpenStudy (ivettef365):

so you have 12 --------- 4x + 2

OpenStudy (ivettef365):

solve as follows: 12 --------------- 4(-1) + 2

OpenStudy (anonymous):

6?

OpenStudy (anonymous):

Since you've been helping me, this is all making so much sense now

OpenStudy (anonymous):

Let f(x) = x + 8 and g(x) = x2 - 6x - 7. Find f(g(2)). -7 -3 10 33 @ivettef365 ??

OpenStudy (anonymous):

When there is the "2" In f(g(2)) what does that mean and how would I go about solving that?

OpenStudy (ivettef365):

on the previews quetions you have 12 /-2 = -6 remember when dividing a positive with negative answer is negative

OpenStudy (ivettef365):

I'm glad I can help

OpenStudy (ivettef365):

on this problem, f(x) = x + 8 and g(x) = x2 - 6x - 7. Find f(g(2)) first step substitute the x for 2 on g(x) g(x) = x^2 - 6x -7 2^2 - 6(2) -7 4 -12-7 4 -19 = -15

OpenStudy (ivettef365):

now you can substitute the x from f(x) for -15, try it to see if you get to answer

OpenStudy (anonymous):

-7 right?

OpenStudy (anonymous):

@ivettef365

OpenStudy (ivettef365):

correct

OpenStudy (anonymous):

Let f(x) = x - 2 and g(x) = x2 - 7x - 9. Find f(g(-1)). -21 -3 3 21

OpenStudy (ivettef365):

now you try doing it and I let you know if is correct. first step substitute the x from g(x) for -1

OpenStudy (anonymous):

I keep running into this problem, (-1)2 - 7(-1) - 9 -2 + 7 - 9 = 4 4 - 2 = 2 and that is none of the possible answers

OpenStudy (anonymous):

0?

OpenStudy (ivettef365):

sorry, my mistake now so (-1)^2 = 1 so you have 1 +7-9 = ??

OpenStudy (anonymous):

-1!

OpenStudy (ivettef365):

correct, now substitute the -1 for the x on f(x)

OpenStudy (anonymous):

-3!

OpenStudy (ivettef365):

correct, good job

OpenStudy (anonymous):

Let f(x) = 2x + 2. Solve f^-1(x) when x = 4. 1 3 4 10 This one is confusing me because of the "f^1" how would I go about doing this and equations like this?

OpenStudy (anonymous):

@ivettef365

OpenStudy (ivettef365):

that is the inverse function

OpenStudy (ivettef365):

sorry, erased it by mistake, here it is again, to find the inverse function f(x) = 2x + 2 you need to write the function as follows: x = 2y + 2 x-2 = 2y x - 2 ------- now substitute the x for the 4 2

OpenStudy (anonymous):

1:)

OpenStudy (anonymous):

Let f(x) = 2x - 6 Solve f-1(x) when x = 2. -5 4 -2 7 Can you show me how I would re-write the inverse function to solve this one?

OpenStudy (anonymous):

I was a little unclear about how to re-write it @ivettef365

OpenStudy (ivettef365):

sorry, I went a little bit fast, because I am at work, and boss was calling me but let me explain, the inverse of a function is just to isolate the x, so in order to do that you get the function f(x) = 2x + 2 and re-write it as x = 2y + 2 ( I substitute the x for y, that way I isolate the y that is on my x value place) x - 2 = 2y (I transfer the 2 to the other side, to transfer you have to change the sign) x - 2 ----- = y (now I divided by 2 on both sides, and the y stays by itself because 2 2y/2 = y) and there you have the inverse

OpenStudy (ivettef365):

now try to do the inverse on this problem you have f(x) = 2x - 6 you write it as x = 2y - 6 (isolate the y)

OpenStudy (anonymous):

x-6/2 = y?

OpenStudy (ivettef365):

remember when you switch a number to the other side need to change the sign

OpenStudy (anonymous):

x+6/2 = y then right?

OpenStudy (ivettef365):

good job, now just substitute the x for 2 that is the easy part

OpenStudy (anonymous):

4!:) Thank you so much for all of your help in teaching me these topics!

OpenStudy (ivettef365):

your welcome :)